OpenStudy (anonymous):

The ratio of an object's weight on Earth to its weight on Neptune is 5 : 7. How much would a person who weighs 150 pounds on Earth weigh on Neptune?

OpenStudy (texaschic101):

you got a proportion here.. 5/7 = 150/x --- (weight of 5 (earth) to 7 (neptune) = 150 (earth) to x (neptune) cross multiply (5)(x) = (7)(150) 5x = 1050 -- divide both sides by 5 x = 210 check.. 5/7 = 150/210 cr0ss multiply (5)(210) = (7)(150) 1050 = 1050 (correct) if the weight on earth is 150, then the weight on neptune is 210

OpenStudy (anonymous):

thankyou

OpenStudy (texaschic101):

anytime :)